1.      Given market demand Qd = 50 - P, and market supply P = Qs + 5

A.      Find the market equilibrium price and quantity

2.      Given utility function U= where PX = 12 Birr, Birr, PY = 4 Birr and the income of the consumer is, M= 240 Birr.

A.      Find the utility maximizing combinations of X and Y.

B.      Calculate marginal rate of substitution of X for Y (MRSX,Y) at equilibrium and interpret your result


1
Expert's answer
2021-08-24T17:06:31-0400

1a

"Qd= 50 - P, P = Qs + 5 -> Qs = P - 5."

The market equilibrium price and quantity are:

"Qd = Qs,\\

50 - P = P - 5,\\

2P = 55,\\

P = \$27.5.

Q = 27.5 - 5 = 22.5 units."

2)

A)



"MU_x=0.5 \frac {y^{0.5}}{x^{0.5}}""MU_y=0.5 \frac {x^{0.5}}{y^{0.5}}""{\frac{MU_x}{p_x}}=\frac {MU_y}{p_y}""x \times p_x+ y \times p_y=M""y=3x""12 \times x+4 \times y=240""x=10, y=30."

D)





"\frac {\partial U}{\partial x \partial y} =\frac {0.25} {(xy)^{0.5}}""MRS x.y = \frac {\partial U} {\partial x \partial y} = \frac {0.25}{(10 \times 30)^{0.5}}=0.015"